OpenStudy (anonymous):

Solve x please show steps 0.00632+0.00632x=0.01-x

OpenStudy (mathstudent55):

Add x to both sides. Subtract 0.00632 from both sides. Then divide both sides by the number multiplying x.

OpenStudy (anonymous):

W/e I did was wrong. 0.00632+0.00632x+x=0.01 0.00632x+x=0.0032 x+x=2.40x10^-5

OpenStudy (anonymous):

@mathstudent55